OpenStudy (anonymous):

Help solving an ODE with Separation of Variables. So I did this with bernouli substitution on an exam and got partial credit because the question said to do this with separation of variables. I feel like this shouldn't be that hard, but I cant for the life of me figure it out. The ODE is: \[y'=y+y ^{-1}\]

OpenStudy (anonymous):

I started with \[\frac{ dy }{ dt } -y = \frac{ 1 }{ y }\]

OpenStudy (anonymous):

Is there some sort of algebra trick I am not seeing? I can't get y by itself to do separation of variables..

OpenStudy (ikram002p):

sound serable to me

OpenStudy (anonymous):

oh man, i think I get it.

OpenStudy (ikram002p):

yeah :) let y and dy in one side and dt in the other side , so what u got ?

OpenStudy (anonymous):

\[(y+y^-1 ) dt\] ug i feel dumb

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=y+y ^{-1}=y+\frac{ 1 }{ y }=\frac{ y^2+1 }{ y }\] separating the variables \[\frac{ y }{ y^2+1 }dy=dx\] integrating

OpenStudy (anonymous):

Ya, ug, thanks a lot, that wasn't bad at all.... ugggggg!

OpenStudy (anonymous):

Thanks for looking guys

OpenStudy (anonymous):

\[\int\limits \frac{ 2y dy }{ y^2+1 }=\int\limits 2 x dx+c\] can you solve further?

OpenStudy (anonymous):

\[\ln \left( y^2+1 \right)=2 \frac{ x^2 }{ 2 }+c\]